OpenStudy (anonymous):

13. Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x^2 +3 = -4x 1, 3 –1, –3 1, –3 1, –3

OpenStudy (anonymous):

Haha im doing this test right now. :P

OpenStudy (anonymous):

haha really? well hey if i get answers ill tell you and if you get answers you you'll tell me?

OpenStudy (anonymous):

Nahh I don't cheat sorry. :)

OpenStudy (anonymous):

aw ok... i usually dont but i really cant afford to get a bad grade now :/ but thats good ur staying honest :)

OpenStudy (anonymous):

Alright. Thanks for the encouragement! :D

OpenStudy (anonymous):

haha welcome! :)

OpenStudy (anonymous):

Ax^2 + Bx = C, or in your case: A = 1, B = 4, C = -3 Quadratic equation = \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Try to solve it, and i'll check ok? :)